Total Manufacturing Cost, Income Statement, Unit Cost, and Selling Price Two inventors, recently organized as Innovation, Inc., consult you regarding a planned new product.They have estimates of the costs of materials, labor, overhead, and other expenses for 2016 but need to know how much to charge for each unit to earn a profit in 2016 equal to 15% of their estimated total long-term investment of $720,000 (ignore income taxes). Their plans indicate that each unit of the new product requires the following: Direct Materia 4 lb. of a material costing $9 per lb. Direct labor 2hrs. of a metal formers time at $19.80 per hr. 0.6 hr. of an assemblers time at $14.40 per hr. Major items of production overhead would be annual rent of $83,628 for a factory building, $51,588 rent for machinery, and $39,060 of indirect material. Other production overhead is estimated to be $419,904 Selling expenses are an estimated 30% of total sales, and non-factory administrative expenses are 20% of total sales. The consensus at Innovation is that during 2016 10,000 units of product should be produced for selling and another 2,000 units should be produced for the next years beginning inventory. Also, an extra 3,000 pounds of material will be purchased as beginning inventory for the next year. Because of the nature of the manufacturing process, all units started must be completed, so work in process inventories are negligible Required a. Incorporate the above data into a schedule of estimated total manufacturing costs and compute the unit production cost for 2016. Do not use negative signs with any of your answersEstimated Total Manufacturing Cost For the Year Ended December 31,2016 Direct material: Beginning materials inventory Cost of material available Less: Total materials used Direct materials used Manufacturing overhead Indirect material Building rent Machinery rent Total manufacturing overhead Total manufacturing costsRound answer to two decimal places. Product cost per unit b. Prepare an estimated income statement that would unt of profit for 2 Income Statement For the Year Ended December 31,2016 Gross profit on sales Operating expenses: C. What unit sales price should Innovation charge for the new product? Check
Answer #1
(a) Computation of total Manufacturing cost and Cost per unit
Estimated Total Manifaturing Cost
For the year ended December 31,2016
Amount in $
Direct materials
Beigining Materials inventory 0
Add: Purchase of materials    ** 459000
Cost of materilas Available 459000
Less: closing Stock of materilas (3000ib *$9) 27000
Total materials used 432000
Direct materilas used 432000
Direct Labour
Metal formars cost =12000*2*$19.80 475200
assembler cost =12000*0.60*$14.40 103680
578880
Manufacturing overhead
Indirect materials 39060
Building rent 83628
Machinery Rent 51588
Other Prodcution over head 419904
Total manufacturing over head 594180
Total manufacturing Costs 1605060
Prodcut Cost per unit = Total manufacturing cost /Production units =$1605060/12000=$133.76
** Purchase cost of materials
Total production = Sales + closing stock =10,000+2000 =12000
Materals required for production =12000*4ib=48000Ib
Total Purcahses =48000ib +3000 ib =51000 ib
Purchases cost of Materials =51000*$9=$459,000
(b) Estimated incoem Statement for the target profit for 2016
Income Statement for the year ended December 31,2016
Amount in $
Total Sales (10,000*$289.12) 2891200
Less: manufacturing cost of goods sold (10,000*$133.76) -1337600
Gross profit 1553600
Less: operating Expenses
Selling Expenses (30% of $2891200) -867360
Non Factory admin Expenses (20% of $2891200) -578240
Net profit 108000
Notes
Target profit =15% *720,000=$108000
Say , Sale price per Unit is 'X', Then total Sales =10000X
Selling and Non factory admin expenses Total 50% of Sales (30%+20%)=10,000X *50% =5000X
Target profit =Gross profit - Selling and Non factory admin expenses
Gross profit = $108000+5000X
Total Sales = Gross profit + manufacturing cost of goods sold =$108,000+5000X +($133.76*10,000)
10,000X =5000X +$14,45,600
5000X =14,45,600
X=14,45,600/5000=$289.12
So, Sale price per Unit to reach the target profit is $289.12